OpenStudy (anonymous):

This is a pretty basic doubt. We say dU (potential energy) = - Force. dx Where it's a dot product. From there, dU = -F.dx *cos(theta) { theta is the angle between F and dx} However, we write F as -dU/dx. Where did the cos(theta) in the dot product vanish?

OpenStudy (ujjwal):

\(\theta\) vanishes only when \(\theta\)=0 That is cos\(\theta\)=1 that is to say when force and displacement are in same direction, cos\(\theta\) vanishes.

OpenStudy (anonymous):

There is a little bit of ambiguity in your question. Writing \[\frac{dU}{dx} = -F \rightarrow dU=-F\cdot dx \] implies that you're working in a one-dimensional system, where vectors are unnecessary. If you're working in multiple spatial dimensions, it would be more appropriate to write \[\vec{\nabla}U = -\vec{F} \rightarrow dU = -\vec{F}\cdot d\vec{r} = -F\cdot dr \cdot \cos(\theta)\] where theta is the angle between the displacement dr and the force F.

OpenStudy (vincent-lyon.fr):

You as using the same notation for F, magnitude of force and Fx, component of force along the x-axis. \(\vec F=-\vec{\nabla}U\) means \(F_x=-\Large\frac{\partial U}{\partial x}\) \(F_y=-\Large\frac{\partial U}{\partial y}\) \(F_z=-\Large\frac{\partial U}{\partial z}\)

OpenStudy (anonymous):

Oh right. It;s the component. That clears it. THanks a lot. :)